# HWC-Loco: A Hierarchical Whole-Body Control Approach to Robust Humanoid Locomotion

### v2

@@ 6 Conclusion
We introduce HWC-Loco, a hierarchical control framework for humanoid robots that incorporates an embedded safety recovery mechanism. This framework has been validated across various locomotion tasks, demonstrating exceptional scalability, robustness, naturalness, and adaptability across diverse tasks and scenarios. Notably, the safety mechanism in HWC-Loco extends beyond locomotion, enabling reliable performance in complex tasks through a dynamic task-safety balance. This ensures robust operation in real-world deployments, positioning HWC-Loco as a foundational solution for safety-critical applications such as industrial automation and assistive robotics. A promising direction for future research is integrating HWC-Loco with upper-body manipulation skills, enabling safety-critical control across a broader range of tasks involving different objects.
Impact Statement
The broader impact of this work lies in its potential to significantly enhance the capabilities of humanoid robots, enabling them to perform complex locomotion tasks safely and effectively across diverse and unpredictable environments.
From an ethical perspective, the development of humanoid robots with advanced locomotion capabilities raises important questions about safety, accountability, and the potential societal implications of widespread adoption. While HWC-Loco prioritizes safety through its robust control mechanisms, ensuring these systems consistently operate within ethical and safety norms is critical to preventing harm, particularly in environments shared with humans. Additionally, as humanoid robots become more capable of replacing human roles in workplaces, concerns about job displacement, economic inequality, and the societal impact of automation must be addressed. Proactive measures, such as workforce reskilling, equitable distribution of benefits, and collaboration between policymakers and researchers, will be necessary to mitigate these risks.
In the long term, HWC-Loco has the potential to significantly improve societal well-being. Humanoid robots equipped with robust locomotion algorithms could assist in tasks that are dangerous or physically demanding for humans, such as search-and-rescue missions, caregiving for the elderly or disabled, and operating in hazardous industrial environments. However, the deployment of such robots must be guided by principles of transparency, fairness, and inclusivity to ensure they serve societal needs without reinforcing biases or creating unintended harms.
Appendix A Additional Training Details
@@ A.1 Motion Retargeting
Our motion data is derived from the CMU Mocap dataset (cmu, ), which exclusively encompasses human locomotion data, featuring a diverse array of walking, jogging, and running styles. In aggregate, we have loaded 318 motion sequences, totaling 3729.18 seconds in duration. Our motion retargeting methodology closely follows the approach outlined in (Cheng et al., 2024). We have meticulously aligned the humanoid robot’s skeleton with that of humans, incorporating adjustments such as rotation, translation, and scaling. Additionally, we map human joints onto the humanoid robot’s structure. Specifically, the human motion data is adapted to the humanoid robot’s framework by mapping local information. Both the Unitree H1 and G1 robots possess shoulder and hip joints that are functionally equivalent to spherical joints. During the retargeting process, the three hip and shoulder joints are treated as a single spherical joint, corresponding to the human’s spherical joint. For one-dimensional joints such as the elbow and torso, the rotation angle is projected onto the corresponding rotation axis of the joint angles.
@@
Table 5: Network Structure
@@ A.2 Implementation Details
Terrain details: Terrains in the training environment simulated by Isaac Gym (Makoviychuk et al., 2021) include flats, obstacles, slopes, and stairs. An example of these terrains is shown in Figure 3.
@@
Low-level Policy Training: For all the low-level policies training, the commands are sampled from ranges: $v_{\text{x}}^{c}\sim[-0.6,2.5]$, $v_{\text{y}}^{c}\sim[-0.6,0.6]$ and $w_{\text{z}}^{c}\sim[-1.0,1.0]$. During the initial training of the humanoid robot on uneven terrain, the robot often remains stationary due to the high difficulty of navigating complex surfaces. To address this, we introduce a terrain curriculum method (Makoviychuk et al., 2021). The training terrain consists of various types, including flat planes, rough surfaces, steps, and slopes. As the robot achieves a goal-tracking performance of 70% of the commanded velocity, the terrain difficulty increases. Conversely, if the goal-tracking performance drops below 40% of the commanded velocity, the terrain becomes easier.
@@
Reward Details: To better facilitate efficient humanoid locomotion, our task reward is defined as (Gu et al., 2024c, d):
@@
To enable the recovery policy to learn stable posture restoration and facilitate smooth policy transitions, we introduce a stand reward. It can be calculated as:
@@
A.3 Saliency Analysis of Observations
We adopt an integrated gradients method similar to (Wang et al., 2024) to assess the importance of different parts of the observation to the policy, as shown in Fig.4. For the recovery policy, ZMP encoding emerges as the most crucial element, followed by the estimated velocity as the second most important factor. Additionally, the robot’s current proprioception and historical proprioception are also significant, while environmental latent has only a slight influence. This is because, for the recovery policy, it is essential to satisfy the ZMP constraints to ensure the robot’s stability and safety. For the high-level policy, ZMP encoding is also the most important component, indicating that it provides vital information for the policy’s decision-making process, thereby enhancing its ability to perceive the robot’s stability. Furthermore, the environmental latent variable proves to be more important compared to the recovery policy, suggesting that environmental factors, such as changes in terrain, can influence the decision-making process.
@@ A.3 Saliency Analysis of Observations
A.4 Extreme Initialization
To simulate extreme scenarios that may occur in real-world deployment, we identify three primary causes of such cases: external disturbances, hardware malfunctions, and malicious velocity inputs. Additionally, we retain domain randomization to enhance the generalization of physical dynamics within the simulation environment. The specific details are outlined as follows:
@@ A.4 Extreme Initialization
4) Domain randomization: Domain randomization is applied to all the training environments. This technique introduces variability in the environment’s parameters, such as friction and gravity. The specifics of the domain randomization are provided in Table 9.
@@
Table 7: Proprioception Noise Parameters
@@
Extreme-Case Analysis: We record nearly 60000 normal state data with randomly sampled commands in diverse terrain settings in 5. We record nearly 10000 extreme cases in the $\mathfrak{P}_{\alpha}^{L}$. Some of the terminal states are visualized in Fig. 6. For the extreme states, the distribution of velocity is close to the command ranges. Roll pitch and their velocity are in a common range, which is near zeros. For the normal states, the distribution of linear velocity overlaps with the command velocity range, with a dense cluster around the high-velocity value of 2 m/s. Regarding yaw velocity, most values fall within the commanded range. However, there are instances of high speed, with some even approaching 10 m/s. This may be due to the robot’s extreme behaviors when encountering unseen situations influenced by external forces and random noise. Regarding the roll-pitch velocity, the pitch velocity shows a higher degree of fluctuation. Similarly, in the roll pitch distribution, for pitch, there is a concentrated distribution at a pitch of 1 m/s, which is one of the terminal conditions. These two figures suggest that the robot is tripped by obstacles during its motion and thus terminates, for instance, when it hits steps or gravel during high-speed motion.
@@
A.5 PPO Algorithm
Table 10: PPO hyperparameters
Hyperparameter Value
Actor Layer [512, 256, 128]
Critic Layer = [512, 256, 128]
Discount Factor 0.99
GAE Parameter 0.95
Epochs per Rollout 5
Minibatch 4
Entropy Coefficient 0.005
Value Loss Coefficient ($\alpha_{1}$) 1.0
Clip Range 0.2
Learning Rate 2.e-4
Environment Steps per Rollout 24
A.6 Discriminator
The discriminator’s hyperparameters are shown in Table. 11.
Table 11: Discriminator hyperparameters
Hyperparameter Value
Discriminator Hidden Layer Dim [1024,512]
Replay Buffer Size 500000
Expert Buffer Size 200000
Expert Fetch Size 512
Learning Batch Size 4096
Learning Rate 2e-5
Gradient Penalty Coefficient 1.0
lamda 1.0
A.7 Deployment Details
@@ Appendix B Experiment
Figure 7: Human-like Behavior
Figure 8: Disturbance: We apply various disturbances to Unitree H1 lasting for one minute. Four representative disturbance segments were selected for analysis. In these segments, a Policy Switching value of 0 corresponds to the goal-tracking policy, while a value of 1 denotes the safety recovery policy. The time is measured in control cycles, with each cycle lasting 0.01 seconds.
Figure 9: Parkour Test: The blue segment represents the goal-tracking policy, while the orange segment signifies the safety recovery policy.

### v3

@@ 7 Conclusion
We introduce HWC-Loco, a hierarchical control framework for humanoid robots that incorporates an embedded safety recovery mechanism. This framework has been validated across various tasks and scenarios, demonstrating exceptional effectiveness, robustness, naturalness, and scalability. Notably, the safety mechanism in HWC-Loco extends beyond locomotion, enabling reliable performance in complex tasks through a dynamic task-safety balance. This ensures robust operation in real-world deployments, positioning HWC-Loco as a foundational solution for safety-critical applications such as industrial automation and assistive robotics. A promising direction for future research is integrating HWC-Loco with loco-manipulation skills, enabling safety-critical control across a broader range of tasks involving different objects.
Appendix A Addtional Details of HWC-Loco
@@ A.1 Motion Retargeting
Our motion data is derived from the CMU Mocap dataset [51], which exclusively encompasses human locomotion data, featuring a diverse array of walking, jogging, and running styles. In aggregate, we have loaded 318 motion sequences, totaling 3729.18 seconds in duration. Our motion retargeting methodology closely follows the approach outlined in [33]. We have meticulously aligned the humanoid robot’s skeleton with that of humans, incorporating adjustments such as rotation, translation, and scaling. Additionally, we map human joints onto the humanoid robot’s structure. Specifically, the human motion data is adapted to the humanoid robot’s framework by mapping local information. Both the Unitree H1 and G1 robots possess shoulder and hip joints that are functionally equivalent to spherical joints. During the retargeting process, the three hip and shoulder joints are treated as a single spherical joint, corresponding to the human’s spherical joint. For one-dimensional joints such as the elbow and torso, the rotation angle is projected onto the corresponding rotation axis of the joint angles.
@@
Table 6: DQN Network Structure
@@ A.2 Implementation Details
Terrain details: Terrains in the training environment simulated by Isaac Gym [63] include flats, obstacles, slopes, and stairs. An example of these terrains is shown in Figure 5.
@@
Low-level Policy Training: For all the low-level policies training, the commands are sampled from ranges: $v_{\text{x}}^{c}\sim[-0.6,2.5]$, $v_{\text{y}}^{c}\sim[-0.6,0.6]$ and $w_{\text{z}}^{c}\sim[-1.0,1.0]$. During the initial training of the humanoid robot on uneven terrain, the robot often remains stationary due to the high difficulty of navigating complex surfaces. To address this, we introduce a terrain curriculum method [63]. The training terrain consists of various types, including flat planes, rough surfaces, steps, and slopes. As the robot achieves a goal-tracking performance of 70% of the commanded velocity, the terrain difficulty increases. Conversely, if the goal-tracking performance drops below 40% of the commanded velocity, the terrain becomes easier.
@@
Reward Details: To better facilitate efficient humanoid locomotion, our task reward is defined as [28, 67]:
@@
The recovery policy shares the same task reward formulation as the goal-tracking policy. However, it uses larger values of the $\sigma_{\text{lin}}^{2}$ and $\sigma_{\text{ang}}^{2}$, which implies a greater tolerance for velocity tracking errors. As a result, this reward term serves as a back-tracking reward for the safety recovery mechanism, encouraging it to return to a stable goal-tracking state. To further promote stable posture restoration and enable smooth transitions back to the goal-tracking policy, we introduce an additional stand reward, defined as:
@@
A.3 PPO
Table 8: PPO hyperparameters
Hyperparameter Value
Actor Layer [512, 256, 128]
Critic Layer [512, 256, 128]
Activation Function ELU
Discount Factor 0.99
GAE Parameter 0.95
Epochs per Rollout 5
Minibatch 4
Entropy Coefficient 0.005
Value Loss Coefficient ($\alpha_{1}$) 1.0
Clip Range 0.2
Learning Rate 2.e-4
Environment Steps per Rollout 24
A.4 VAE
The VAE Network consists of an encoder and a decoder. The encoder takes historical observations as input and outputs latent representation and estimation of priviledged information. The decoder takes latent as input and outputs the reconstruction of the next observation. The structure of VAE is shown in Figure 6
Figure 6: VAE Structure
Table 9: VAE hyperparameters
Hyperparameter Value
Encoder Layer [256, 128]
Decoder Layer [256, 128, 64]
Activation Function ELU
KL Coefficient 1.0
Learning Rate 1.e-4
A.5 Discriminator
The discriminator’s hyperparameters are shown in Table. 10.
Table 10: Discriminator hyperparameters
Hyperparameter Value
Discriminator Hidden Layer Dim [1024,512]
Replay Buffer Size 500000
Expert Buffer Size 200000
Expert Fetch Size 512
Learning Batch Size 4096
Learning Rate 2e-5
Gradient Penalty Coefficient 1.0
lamda 1.0
A.6 Saliency Analysis of Observations
We adopt an integrated gradients method similar to [70] to assess the importance of different parts of the observation to the policy, as shown in Fig.7.
For the recovery policy, the Zero Moment Point (ZMP) encoding proves to be the most critical factor, followed by the estimated velocity. The robot’s current and historical proprioceptive states also contribute substantially to the policy’s decision-making process. In contrast, the environmental latent representation exhibits only a marginal effect. This hierarchy of importance reflects the necessity of strictly adhering to ZMP constraints to maintain dynamic stability and ensure the safety of the robot during recovery.
For the high-level policy, the Zero Moment Point (ZMP) encoding remains the most influential component, indicating that it provides critical information for decision-making and substantially enhances the policy’s awareness of the robot’s stability. This observation implies that the high-level policy prioritizes stability-relevant features over task-irrelevant ones, which is advantageous for promoting generalization across diverse scenarios. Moreover, the environmental latent variable exhibits greater importance compared to that in the recovery policy, suggesting that high-level decisions are more sensitive to environmental factors such as terrain variations.
@@ A.6 Saliency Analysis of Observations
A.7 Extreme Initialization
To simulate extreme scenarios that may occur in real-world deployment, we identify three primary causes of such cases: external disturbances, hardware malfunctions, and malicious commands. Additionally, we retain domain randomization to enhance the generalization of physical dynamics within the simulation environment. The specific details are outlined as follows:
To enhance robustness and generalization in simulation, we simulate both extreme deployment scenarios and apply domain randomization. The extreme scenarios include external disturbances, sensor and actuator failures, and adversarial commands, while domain randomization targets physical variability in the environment.
@@ A.7 Extreme Initialization
4) Domain randomization (for environment variability): Domain randomization is applied to all the training environments. This technique introduces variability in the environment’s parameters, such as friction and gravity. The specifics of the domain randomization are provided in Table 13.
@@
Table 11: Intense Noise Parameters
@@
Extreme-Case Analysis: We record nearly 60000 normal state data with randomly sampled commands in diverse terrain settings in 8. We record nearly 10000 extreme cases in the $\mathfrak{P}_{\alpha}^{L}$. Some of the terminal states are visualized in Fig. 9. For the extreme states, the distribution of velocity is close to the command ranges. Roll pitch and their velocity are in a common range, which is near zeros. For the normal states, the distribution of linear velocity overlaps with the command velocity range, with a dense cluster around the high-velocity value of 2 m/s. Regarding yaw velocity, most values fall within the commanded range. However, there are instances of high speed, with some even approaching 10 m/s. This may be due to the robot’s extreme behaviors when encountering unseen situations influenced by external forces and random noise. Regarding the roll-pitch velocity, the pitch velocity shows a higher degree of fluctuation. Similarly, in the roll pitch distribution, for pitch, there is a concentrated distribution at a pitch of 1 m/s, which is one of the terminal conditions. These two figures suggest that the robot is tripped by obstacles during its motion and thus terminates, for instance, when it hits steps or gravel during high-speed motion.
@@
A.8 History Length Experiments
We investigate the impact of observation history length on HWC-Loco’s performance. Setting $H=10$ achieves the best overall trade-off across the three metrics. While increasing $H$ beyond 10 slightly improves the success rate, it leads to a noticeable decline in goal-tracking accuracy and human-likeness. This suggests that excessively long history may hinder the effective use of privileged information, resulting in more conservative and less adaptive behaviors.
Table 14: Performance comparison with different history lengths ($H$)
History Length Success Rate $\uparrow$ Goal-tracking $\uparrow$ Human-like $\downarrow$
1 $95.13\pm 0.51$ $1.07\pm 0.00$ $3.26\pm 0.06$
10 $96.73\pm 0.40$ $\mathbf{1.10\pm 0.00}$ $\mathbf{3.18\pm 0.01}$
20 $96.53\pm 0.40$ $1.08\pm 0.00$ $3.18\pm 0.02$
40 $\mathbf{97.38\pm 0.33}$ $1.02\pm 0.00$ $3.21\pm 0.02$
A.9 Training Cost
All policies are trained on an NVIDIA RTX 4090 GPU with 4,096 parallel environments. Our framework involves training three policies. The goal-tracking policy converges within approximately 10000 iterations (nearly 10 hours), while the recovery policy requires around 8000 iterations to converge (nearly 8 hours). The high-level policy demonstrates faster convergence, requiring approximately 6000 iterations (under 6 hours).
A.10 Deployment Details
@@
During deployment, inference operates through a hierarchical structure involving both high-level and low-level policies. To meet the strict latency and resource constraints of onboard systems, the policies are optimized for lightweight designs and exported JIT format. The high-level policy model size is below 2MB, and the low-level policies remain under 3MB. Those designs ensures that real-time inference at 100 Hz is feasible on the onboard computer of the Unitree H1 robot.
@@ Appendix B Experiment
B.1 H1 Experiments
Effectiveness Tests in the Real World. Figure 10 demonstrates the robot’s ability to traverse a multi-terrain course consisting of 15 cm steps, 20° slopes, and wooden boards at a speed of 0.3 m/s. Figure 11 shows the ability to stability in real-world outdoor locomotion, including flat ground, grassy surfaces, and slopes.
Figure 10: Climb Stairs Test. The blue segments indicate the activation of the goal-tracking policy, while the orange segments correspond to the safety recovery policy.
(a) Walking on flat terrain.
(b) Traversing grassy surfaces.
(c) Climbing a slope.
(d) Descending a slope.
Figure 11: Outdoor Deployment across Diverse Terrains: (a) flat ground, (b) grassy surface, (c) upward slope, and (d) downward slope.
Robustness Tests in the Real World. As shown in Figure 12, we apply various disturbances to Unitree H1 lasting for one minute. Four representative disturbance segments were selected for analysis. The time is measured in control cycles, with each cycle lasting 0.01 seconds. Furthermore, we conduct an outdoor disturbance test by applying pushes, pulls, and kicks to the robot, as shown in Figure 13. The robot waves its arms and adjusts its gaits to maintain balance. Notably, the high control frequency (100 Hz) ensures smooth transitions between policies, and isolated activations of the recovery policy have minimal impact on overall stability.
Figure 12: Disturbances in Realistic Deployment: Policy Switching value of 0 corresponds to the goal-tracking policy, while a value of 1 denotes the safety recovery policy.
Figure 13: Robustness in Outdoor Settings: The robot responds to external disturbances in an outdoor environment by waving its arms and adjusting its gaits to maintain balance.
Case Study on Motion Tracking. We visualize the robot’s performance in tracking a punching motion. During the execution, the robot experiences an impulse impact. In response, it dynamically activates the recovery policy, switching to more conservative and safer actions. Specifically, it lowers its body, adjusts its gait, and stabilizes itself to satisfy ZMP constraints, rather than strictly following the original target motion. As a result, it generates a backtracking motion to maintain balance and ensure robustness.
Figure 14: Robustness in Motion Tracking: The green dots represent the target keypoints, the blue dots indicate the keypoints corresponding to the DoFs, and the red dot denotes the ZMP.
B.2 G1 Experiments
Table 16: Comparison of Humanoid Configurations
Parameter H1 G1
Height 178 cm 127 cm
Mass 47 kg 35 kg
DoFs 19 23
G1 outperforms H1 across all three metrics (success rate, goal-tracking accuracy, and human-likeness) primarily due to the following factors:
  • Size and Mass: As shown in Table 16, G1’s smaller size and lighter mass contribute to lower inertial loads and reduced torque demands. This facilitates greater stability, especially during dynamic motions, and enables faster corrective actions in response to disturbances.
  • Degrees of Freedom (DoFs): With 23 DoFs compared to H1’s 19, G1 offers enhanced joint-level flexibility. The additional DoFs, particularly in the upper body and feet, allow for better whole-body coordination. This increased redundancy leads to improved disturbance rejection and more human-like behavior during complex locomotion tasks.
Notably, as illustrated in Figure 15, G1 is capable of achieving straight-knee walking, which remains a significant challenge for H1 due to its larger size and reduced agility. The additional DoFs in G1 enable finer control of limb trajectories, which is critical for maintaining balance during such biomechanically demanding gaits.
Figure 15: G1 walk with straight knees